OpenStudy (anonymous):

PLEAASSEEE Explain, really confused In a laboratory experiment, pure lead metal reacted with excess sulfur to produce a lead sulfide compound. The following data was collected: Mass of empty evaporating dish: 25.000 g Mass of evaporating dish and lead metal: 26.927 g Mass of evaporating dish and lead sulfide: 27.485 g Solve for the empirical formula of the lead sulfide compound. Be sure to show, or explain, all of your calculations.

OpenStudy (anonymous):

okay,from these information, amount of lead=26.927-25.000gm=1.927gm, amount of lead sulfide formed=27.485-25.000gm=2.485gm, from here,amount of sulfur is=2.485-1.927=.558gm, now here says that excess of sulfur is taken,so here minimum amount of lead sulfide formed will be determined by the amount of lead only.so limiting agent is lead. so, now we have take the amount of mole is used for forming the compound, for lead it will be=1.927/206.9=.009 for sulfur it will be=.558/64=.008 now using simplest ratio 1 mole of lead is reacted with 1 mole of sulfur to produce 1 mole of lead sulfide,so the empirical formula will be PbS

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

get it??

OpenStudy (anonymous):

good